Project

General

Profile

Actions

Ops/Comm Scheduling meeting

Location : CR21 in R1 @ 15:30 pm on October 4th

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

Hall work

  • Shutdown activities : 28 October - 15 November
    • Mop-up of summer list and other issues?
    • SSD cryocooler system refurb (Nov 9)
    • Possible test of new Linde control PC
    • Conventional magnets: Earth leakage tests at magnets (inc. synch and DSA)

Operations

  • MOM report
  • Detectors status (1 minute/1 slide)
    • Melissa - Tracker
    • Mariyan - KL
    • Scott - TOF
    • Francois - EMR
    • Lucien - Ckov
  • Run Plan
    • Owl shifts

Notes

Chris Rogers
Victoria
Henry
Craig
Paolo
Victoria
Mark
Josef
Milorad
Ed
Durga
Ken
Colin
Paul
Pierrick
Jan
Alan

DC Report
  • Air con in RR2 fixed; valve was broken; this was repaired, now one is running, one is on standby; they will alternate once a week between standby and running
  • Water; nominally complete, controls system is working, can change temperature from GUI in MLCR; dont have roof system sorted, limited by cooling of water that circulates into the hall. Appear to have cooling capacity in the roof but it is not transferring down. Discussion with controls person tomorrow; using 300-400 kW power budget which is close to capacity. Henry did a test run on Thursday evening and now another run with Henry settings is ongoing. Change in temperature seems to be worse today than Thursday. Net heat load is comparable to cooling channel settings. Colin: nothing has changed in principle since Thursday. Henry: note diurnal variation of ~ 2 degrees C. Margin should get better at night. We agree to take data at lower conventional magnet currents and try 140 MeV/c.
    - Next aim to establish 140 MeV/c
Decay solenoid
  • Full of He and we will power test tomorrow; question will we need a spare for the air end that failed? No there was a suspect seal. Need access to the power supply tomorrow.
Tracker
  • Milorad has arrived today with two spare boards; both boards powered on fine. Trying to install firmware; needs software install which is not yet complete. Plan to get firmware on ASAP, for install in the hall tomorrrow.
Diffuser
  • Mike Dawson coming up from Oxford tomorrow; visual inspection and local repair of diffuser control or ship to Oxford.
Maintenance day list
  • Add Test decay solenoid
  • VLSB replacement on tracker
  • Melissa - hall probe work
  • Pierrick controls work - decided not to do development work on the power supplies
    Hall open or closed?
  • hall should be open
    Magnets powered?
  • Not compatible with chiller work
  • Can do a field survey to verify field level on the roof; but the
    Should we power cycle the magnets? Josef expresses no strong opinion.
So the plan:
  • Confirm power down magnets tonight. Power down at 9 pm tonight. Pierrick and Josef will be supervise ramp down. Rogers will supervise going to free access.
  • Lock up and handover Wednesday evening for straight track data.
  • Ramp up Thursday morning, noting requirement for water temperature.
12th October maintenance day
  • Only item so far is TOF work in DSA. We need health physics before entry to DSA. TOF folks need a plan - ACTION on DC
SSD Cryocooler refurb
  • Colin: Need clarity about what is required Pierrick: internal cleaning of the Helium system; and clean up of the vacuum space. Believe plating on the coldmass and to clean it out we need to do a partial warm up to pump it out. Changing of the Helium in the compressor circuits can be done individually on each compressor. Warm up of coldmass - what is the risk vs benefit? Before found lots of release, have to actively maintain cooling. Colin wants evidence - 1. that we can solve the problem; 2. that we have a problem; 3. that we can prevent a run away situation. Need some careful risk assessment. Mike: how will you rid yourself of plated Nitrogen?
MOM report
  • ISIS was off Tuesday through Thursday following failed main magnet; Friday and Saturday we took TOF calibration work. Sunday we were off due to target hardware failure. Monday afternoon/evening ISIS was off. Upcoming week we will do the magnet scans and more scattering data. Monday, we expect the LiH install. Note that we have no magnet babysitters. Steve will beg for shifters during the plenary; Rogers will beg for shifters during analysis parallel.

Updated by Nebrensky, Henry over 6 years ago ยท 19 revisions